OpenStudy (anonymous):

A 1.0 kg weight suspended from a spring is pulled to 0.25 m below its equilibrium point. If the spring has a spring constant (K) of 50.0 N/m, at what rate will the mass accelerate when it is released? 6.5 m/s^2 2.7 m/s^2 26 m/s^2 13 m/s^2 I think it's A, but I'm unsure.

OpenStudy (shamim):

U know F=kx=?

OpenStudy (shamim):

Given k nd x

OpenStudy (shamim):

Then use F=ma a=?

OpenStudy (kaptain_mittens):

So do F=kx as shamim said and you end up with: F=.25(50), then after getting F you do F=1kg(a)

OpenStudy (anonymous):

Okay @kaptain_mittens so I get f=12.5 12.5=1kg(a)?

OpenStudy (shamim):

Tell us a=?

OpenStudy (anonymous):

I think that's why I'm confused :( @shamim

OpenStudy (anonymous):

The force of a spring can be given as: \[\huge \text{F}_\text{spring}=k \Delta x\]Given your scenario, we have\[\huge \text{F}_\text{spring}=(0.25)(50)=12.5\]However, given that it's a force, any force will follow Newton's 2nd Law, which is\[\huge \text{F}=ma\]Using our spring force that we calculated earlier, we can see that\[\huge 12.5=(1)(a) \rightarrow a=12.5 \frac{\text{m}}{\text{s}^2}\]